# Video Deepfake Abuse: How Company Choices Predictably Shape Misuse Patterns

### v1

@@ 4.1. How can Developers Reduce Risks?
Existing literature on AI safeguards and risk management establishes that several techniques can significantly mitigate harmful uses of open-weight AI models (casper2025open). Here, in line with casper2025open, we focus on risk mitigation techniques for open-weight models that cannot be easily circumvented by users. [^8] In addition to risk mitigation techniques that we present in Section 4.1, data provenance methods also serve an important role. While these techniques do not directly mitigate risks, they facilitate a broader understanding of the ecosystem. See bengio2025international and casper2025open. For example, while external safety filters can be valuable to deploy alongside open-weight models (e.g. agarwal2025cosmos), they can be trivially disabled and are empirically weak safeguards. Crucially, comprehensive AIG-NCII risk mitigation strategies for open models must account for how fine-tuned derivatives of open-weight models, rather than the original models themselves, are empirically a major contributor to misuse (simonovich2024nytheon; casper2025open). For example, CivitAI hosts thousands of fine-tuned low-rank adapters designed to make video generation models more consistent, dynamic, or specialized in their ability to produce NSFW content.
@@ 4.1.1. Training Data Curation
Current evidence suggests that data filtering is a useful intervention for mitigating harmful content generation capabilities. Filtering content from training data related to harmful topics helps developers train models with minimal knowledge of such content (maini2025safety; o2025deep; nichol2021glide). While the relationship between training data filtering, model capabilities, and resistance to harmful fine-tuning is not yet fully understood in diffusion models, filtering NSFW content from training data is recognized as a key defense against AIG-NCII (NIST_ID012_2024; NIST_ID012_2024). A model trained on NSFW-filtered data can always be taught to generate NSFW content after enough fine-tuning, but removing NSFW content from training data substantially increases barriers to misuse by requiring users to fine-tune new NSFW capabilities into a model before it can be misused (o2025deep; nichol2021glide). Dataset curation can also be an important point of harm remediation in machine learning research, as some datasets used in open science contain non-consensually obtained datasets of NSFW content (Cintaqia2025Stop).
@@
Stable Diffusion 2.0 provides empirical evidence for the effectiveness of filtering NSFW data as a safeguard. Stability AI released Stable Diffusion 2.0 in November 2022, reporting that “Stable Diffusion 2.0 delivers several big improvements and features versus the original V1 release” (StableDiffV2Release2022). Stable Diffusion 2.0 was trained on “an aesthetic subset of the LAION-5B dataset…further filtered to remove adult content using LAION’s NSFW filter” (StableDiffV2Release2022). Subsequently, the NSFW content generation community has empirically found Stable Diffusion 2.0 to be unhelpful for generating NSFW content. For example, we identified various (SFW) Reddit posts reporting “We all know by now that SD 2.0 is pretty poor for celebrities, artists, and NSFW content,” that Stable Diffusion 2.0 was “abandoned because it can not generate any NSFW content,” and “Well if you’re an overly sensitive prude who has anxiety attacks if you see some [redacted], SD 2.0 is great because it’s so heavily censored.” We confirmed this using analysis of search hits on CivitAI (Figure 3). Across three search terms (“nsfw”, “nude”, and “naked”), Stable Diffusion 1.x models and derivatives had dramatically more NSFW content than Stable Diffusion 2.x models: 179,345+ total hits compared to only 128, resulting in a ratio of 1,401×. [^10] The “nsfw” search for Stable Diffusion 1.x exceeded CivitAI’s display limit of 100,000 results, so the actual difference is likely substantially larger. Meanwhile, the ratio of NSFW to NSFW fine-tuned models is over twice as large for Stable Diffusion 1.x.
@@ 4.1.2. Fine-Tuning, Unlearning, and Anti-Tampering
Post-training safeguards can significantly increase barriers to misuse. After a diffusion model is trained, post-training safeguards can offer a second line of defense against misuse of NSFW content generation capabilities. In particular, developers can utilize “machine unlearning” methods to suppress NSFW-generation capabilities. In recent years, researchers have introduced and tested a number of approaches to unlearning specific capabilities from diffusion models, including NSFW content (zhang2024defensive; fuchi2024erasing; zhu2024choose; wu2025erasing; zhang2025concept; wu2025unlearning; lu2025concepts). To date, the bulk of research on these methods has focused on image diffusion models, but the same algorithms are equally applicable to video diffusion models (e.g., liu2024unlearning). While useful, most current unlearning methods are brittle and can be partially circumvented by adversarial prompts or adversarial few-shot fine-tuning (sharma2024unlearning; suriyakumar2024unstable; zhang2024generate; george2025illusion; yu2025forgetme). However, there is also emerging research on diffusion model unlearning methods that resist relearning (gao2024meta; li2025towards; abdalla2025gift). For example, abdalla2025gift introduced an unlearning method that caused diffusion models to resist up to 2,500 steps of fine-tuning on NSFW content while also preserving the model’s ability to learn from benign images. Overall, research is still ongoing to develop more practical and relearning-resistant unlearning techniques for video diffusion models. Nonetheless, existing post-training methods for suppressing NSFW generation capabilities offer a barrier to misuse without significantly affecting the SFW capabilities of a diffusion model.
@@ 4.1.3. Evaluations
Evaluating open-weight models for misuse potential, particularly under fine-tuning threats, is needed to meaningfully assess real-world misuse risks. Evaluations (both internal and external) of frontier AI systems’ capabilities are increasingly recognized as a key pillar of AI risk management frameworks. Safety evaluations of frontier systems are most commonly conducted using input-output access in which evaluators give models queries and analyze their responses. For closed weight models, input-output evaluations emulate the real-world misuse threats that models will be exposed to. However, for open-weight models, input-output evaluations do not fully account for the risks posed by simple downstream modifications (casper2024black; che2025model; casper2025open). While it is intuitive that evaluating the safety of open-weight models requires adversarial fine-tuning, there is limited precedent for open-weight models to be evaluated under harmful fine-tuning attacks. For example, while some of the developers we analyzed in Section 3 report on evaluations for NSFW capabilities (see Table 1), zero mention evaluations under fine-tuning.
@@ 4.1.4. Staged Deployment
Staged release strategies enable monitoring and patches prior to an open-weight release. Model deployment is not a binary between fully closed and fully open. Different deployment strategies offer different trade-offs between openness and security. When a company has the resources to implement them, these options allow model developers to gradually increase access to a model before choosing to fully release it with open weights (solaiman2023gradient). For example, a developer could first provide API query access to a model to study its input-output uses. Second, they could offer access to a fine-tuning API to study risks from fine-tuning. Third, they could release only some layers of the model and enable access to the others through a collaborative inference API (ren2023survey). These strategies allow developers to monitor how real users interact with their systems, identify emerging misuse patterns, and refine safety measures before a full open release (casper2025open). For example, observations following a staged deployment could help the developer identify a need for additional mitigations, such as unlearning (Section 4.1.2).
4.2. Developers Rarely Report on Mitigations
Most open-weight model developers do not report on efforts to safeguard their models against AIG-NCII misuse. Above in Section 3.2, we observe that some models, including Wan2.1, stable-video-diffusion, HunyuanVideo, and LTX-Video are disproportionately used for NSFW content. To better understand why this might be the case, we analyze technical reports from developers for mentions of the mitigation strategies discussed above in Section 4.1.111111Our analysis was limited to reports available online in English. However, some of the developers in Table 1 are based in countries where English is not the official language. Further details may be available in non-English languages. We summarize results for the same ten models as in Section 3.2 in Table 1. With the exception of Cosmos (agarwal2025cosmos), developers consistently report very little about AIG-NCII risks or risk mitigations. And even when information is presented, it is often with very little detail.
@@ 4.3. How Can Model Distribution Platforms Reduce Risks?
Model distribution platforms are key to the accessibility of harmful model capabilities. Thus far, we have primarily focused on video model developers because development is a key point of influence and leverage in the AI ecosystem (williams2025regulating). However, while developers are responsible for creating models, distribution platforms are principally responsible for their widespread accessibility (gorwa2024moderating).
@@
Distribution platforms can amplify misuse by hosting and distributing models designed for AIG-NCII. Many early image-generation tools for of generating AIG-NCII were hosted and distributed on GitHub, one of the most popular developer platforms in the world. In 2023, the National Center on Sexual Exploitation reported that GitHub had been hosting open-source repositories that enable AIG-NCII generation. This included DeepFaceLab, which “contains direct links to the most prolific sexual deepfake abuse website in the United States: Mr.DeepFakes,” as well as repositories for DeepNude and Unstable Diffusion (patrick_trueman_esq_github_2023).
@@
In the past several years, generative AI model distribution has increasingly shifted to CivitAI and Hugging Face. In May 2025, following new regulations and pressure from payment processors, CivitAI banned models designed to recreate real individuals’ likenesses, which, according to 404 Media, dealt a “major blow to the nonconsensual AI porn ecosystem” (civitai404media; civitai404media1). However, the deepfake community subsequently organized a coordinated effort to archive over 5,000 of these models and re-upload them to Hugging Face (maiberg2025). Recently, Hugging Face has been found to host models specifically designed for producing AIG-NCII of celebrities, including individuals who were minors when source material was created (hashim2025; maiberg2025). These models violated Hugging Face’s stated content policy prohibiting “sexual content created without explicit consent” and “underage nudity or any sexual content involving minors” (hfcontentpolicy2025). Nonetheless, maiberg2025 found that Hugging Face was slow to take these models down, allowing some to remain accessible for nearly five months despite multiple requests. Even after removal events, some violating models remained accessible (hashim2025). This pattern illustrates how a lack of proactive enforcement of anti-AIG-NCII policies on model distribution platforms can make harmful models significantly more widely accessible.
@@ 5.1. Are AIG-NCII Harms from Open Video Generators Unmitigatable?
The “cat-out-of-the-bag” argument overlooks harm reduction principles and incorrectly characterizes technology adoption dynamics. The ‘cat-out-of-the-bag’ reasoning is not in line with established approaches to risk management and misidentifies how technology proliferates. Mitigations matter. For example, digital piracy provides one clear parallel: despite the theoretical availability of pirated content online, coordinated enforcement efforts by search engines, authorities, and companies to reduce the accessibility of pirated content online have had a measurable impact. For example, smith2023piracy concluded that removing links to pirated content from search results increases legal content sales by 11-14%, and (danaher2017) concluded that search engines that demote piracy sites in results shift user behavior toward legal consumption. Even though determined users can still find pirated content with sufficient effort, increasing barriers to accessing it meaningfully reduces piracy by increasing search costs and directing casual users toward legitimate alternatives. Reducing the accessibility of harmful generative AI capabilities may be particularly helpful for reducing AIG-NCII given that a large number of perpetrators are minors themselves (alexander2025deepfake).
Risk mitigation operates through aggregate effects, creating systemic benefits even when circumvention remains possible. Safety measures implemented by major platforms and developers create systemic benefits, including reducing total explicit content volume by avoiding accidental harms, preventing unsophisticated misuse, deterring costly circumventions of safeguards, preventing normalization by keeping harmful applications marginalized rather than mainstream, and enabling legal frameworks that can target deliberate safety circumvention (ec2019).
Technology adoption follows network effects that amplify early choices made by major actors. The first widely available implementations establish user expectations, development priorities, and community norms that influence subsequent development. For example, the LAION dataset was introduced in March 2022 (laion2022laion5b), was found to contain CSAM in December 2023 (thiel2023), and was modified to mitigate CSAM content in August 2024 (laion2024relaion5b). Nonetheless, prior to August 2024, the original LAION dataset was widely used to train generative systems, including Stable Diffusion models (Salvaggio2024), that are still commonly used today. When major developers and platforms implement robust safety measures, they also create incentives for compatible approaches across the ecosystem. Conversely, early releases without safety measures normalize applications and create communities focused on circumvention rather than beneficial use. As such, the “cat-out-of-the-bag” argument offers a pretext to engage in a race to the bottom.

### v2

@@ 4.1. How can Developers Reduce Risks?
Existing literature on AI safeguards and risk management establishes that several techniques can significantly mitigate harmful uses of open-weight AI models (Casper et al., 2025). Here, in line with Casper et al. (2025), we focus on risk mitigation techniques for open-weight models that cannot be easily circumvented by users. [^8] In addition to risk mitigation techniques that we present in Section 4.1, data provenance methods also serve an important role. While these techniques do not directly mitigate risks, they facilitate a broader understanding of the ecosystem. See Bengio et al. (2025) and Casper et al. (2025). For example, while external safety filters can be valuable to deploy alongside open-weight models (e.g. Agarwal et al., 2025), they can be trivially disabled and are empirically weak safeguards. Crucially, comprehensive AIG-NCII risk mitigation strategies for open models must account for how fine-tuned derivatives of open-weight models, rather than the original models themselves, are empirically a major contributor to misuse (Simonovich, 2024; Casper et al., 2025). For example, CivitAI hosts thousands of fine-tuned low-rank adapters designed to make video generation models more consistent, dynamic, or specialized in their ability to produce NSFW content.
@@ 4.1.1. Training Data Curation
Current evidence suggests that data filtering is a useful intervention for mitigating harmful content generation capabilities. Filtering content from training data related to harmful topics helps developers train models with minimal knowledge of such content (Maini et al., 2025; O’Brien et al., 2025; Nichol et al., 2021). While the relationship between training data filtering, model capabilities, and resistance to harmful fine-tuning is not yet fully understood in diffusion models, filtering NSFW content from training data is recognized as a key defense against AIG-NCII (Thorn and ATIH, 2024, 2024). A model trained on NSFW-filtered data can always be taught to generate NSFW content after enough fine-tuning, but removing NSFW content from training data substantially increases barriers to misuse by requiring users to fine-tune new NSFW capabilities into a model before it can be misused (O’Brien et al., 2025; Nichol et al., 2021). Dataset curation can also be an important point of harm remediation in machine learning research, as some datasets used in open science contain non-consensually obtained datasets of NSFW content (Cintaqia et al., 2025).
@@
Stable Diffusion 2.0 provides empirical evidence for the effectiveness of filtering NSFW data as a safeguard. Stability AI released Stable Diffusion 2.0 in November 2022, reporting that “Stable Diffusion 2.0 delivers several big improvements and features versus the original V1 release” (StabilityAI, 2022). Stable Diffusion 2.0 was trained on “an aesthetic subset of the LAION-5B dataset…further filtered to remove adult content using LAION’s NSFW filter” (StabilityAI, 2022). Subsequently, the NSFW content generation community has empirically found Stable Diffusion 2.0 to be unhelpful for generating NSFW content. For example, we identified various (SFW) Reddit posts reporting “We all know by now that SD 2.0 is pretty poor for celebrities, artists, and NSFW content,” that Stable Diffusion 2.0 was “abandoned because it can not generate any NSFW content,” and “Well if you’re an overly sensitive prude who has anxiety attacks if you see some [redacted], SD 2.0 is great because it’s so heavily censored.” We confirmed this using analysis of search hits on CivitAI (Figure 3). Across three search terms (“nsfw”, “nude”, and “naked”), Stable Diffusion 1.x models and derivatives had dramatically more NSFW content than Stable Diffusion 2.x models: 179,345+ total hits compared to only 128, resulting in a ratio of 1,401×. [^10] The “nsfw” search for Stable Diffusion 1.x exceeded CivitAI’s display limit of 100,000 results, so the actual difference is likely substantially larger. Meanwhile, the ratio of NSFW to NSFW fine-tuned models is over twice as large for Stable Diffusion 1.x.
@@ 4.1.2. Fine-Tuning, Unlearning, and Anti-Tampering
Post-training safeguards can significantly increase barriers to misuse. After a diffusion model is trained, post-training safeguards can offer a second line of defense against misuse of NSFW content generation capabilities. In particular, developers can utilize “machine unlearning” methods to suppress NSFW-generation capabilities. In recent years, researchers have introduced and tested a number of approaches to unlearning specific capabilities from diffusion models, including NSFW content (Zhang et al., 2024a; Fuchi and Takagi, 2024; Zhu et al., 2024; Wu et al., 2025a; Zhang et al., 2025; Wu et al., 2025b; Lu et al., 2025). To date, the bulk of research on these methods has focused on image diffusion models, but the same algorithms are equally applicable to video diffusion models (e.g., Liu and Tan, 2024; Xu et al., 2025; Ye et al., 2025). While useful, most current unlearning methods are brittle and can be partially circumvented by adversarial prompts or adversarial few-shot fine-tuning (Sharma et al., 2024; Suriyakumar et al., 2024; Zhang et al., 2024b; George et al., 2025; Yu et al., 2025). However, there is also emerging research on diffusion model unlearning methods that resist relearning (Gao et al., 2024; Li et al., 2025; Abdalla et al., 2025). For example, Abdalla et al. (2025) introduced an unlearning method that caused diffusion models to resist up to 2,500 steps of fine-tuning on NSFW content while also preserving the model’s ability to learn from benign images. Overall, research is still ongoing to develop more practical and relearning-resistant unlearning techniques for video diffusion models. Nonetheless, existing post-training methods for suppressing NSFW generation capabilities offer a barrier to misuse without significantly affecting the SFW capabilities of a diffusion model.
@@ 4.1.3. Evaluations
Evaluating open-weight models for misuse potential, particularly under fine-tuning threats, is needed to meaningfully assess real-world misuse risks. Evaluations (both internal and external) of frontier AI systems’ capabilities are increasingly recognized as a key pillar of AI risk management frameworks. Safety evaluations of frontier systems are most commonly conducted using input-output access in which evaluators give models queries and analyze their responses. For closed weight models, input-output evaluations emulate the real-world misuse threats that models will be exposed to. However, for open-weight models, input-output evaluations do not fully account for the risks posed by simple downstream modifications (Casper et al., 2024; Che et al., 2025; Casper et al., 2025). While it is intuitive that evaluating the safety of open-weight models requires adversarial fine-tuning, there is limited precedent for open-weight models to be evaluated under harmful fine-tuning attacks. For example, while some of the developers we analyzed in Section 3 report on evaluations for NSFW capabilities (see Table 1), zero mention evaluations under fine-tuning.
@@ 4.1.4. Staged Deployment
Staged release strategies enable monitoring and patches prior to an open-weight release. Model deployment is not a binary between fully closed and fully open. Different deployment strategies offer different trade-offs between openness and security. When a company has the resources to implement them, these options allow model developers to gradually increase access to a model before choosing to fully release it with open weights (Solaiman, 2023). For example, a developer could first provide API query access to a model to study its input-output uses. Second, they could offer access to a fine-tuning API to study risks from fine-tuning. Third, they could release only some layers of the model and enable access to the others through a collaborative inference API (Ren et al., 2023). These strategies allow developers to monitor how real users interact with their systems, identify emerging misuse patterns, and refine safety measures before a full open release (Casper et al., 2025). For example, observations following a staged deployment could help the developer identify a need for additional mitigations, such as unlearning (Section 4.1.2).
@@
4.2. Developers Rarely Report on Mitigations
Most open-weight model developers do not report on efforts to safeguard their models against AIG-NCII misuse. Above in Section 3.2, we observe that some models, including Wan2.1, stable-video-diffusion, HunyuanVideo, and LTX-Video are disproportionately used for NSFW content. To better understand why this might be the case, we analyze technical reports from developers for mentions of the mitigation strategies discussed above in Section 4.1.111111Our analysis was limited to reports available online in English. However, some of the developers in Table 1 are based in countries where English is not the official language. Further details may be available in non-English languages. We summarize results for the same ten models as in Section 3.2 in Table 1. With the exception of Cosmos (Agarwal et al., 2025), developers consistently report very little about AIG-NCII risks or risk mitigations. And even when information is presented, it is often with very little detail.
@@ 4.3. How Can Model Distribution Platforms Reduce Risks?
Model distribution platforms are key to the accessibility of harmful model capabilities. Thus far, we have primarily focused on video model developers because development is a key point of influence and leverage in the AI ecosystem (Williams et al., 2025). However, while developers are responsible for creating models, distribution platforms are principally responsible for their widespread accessibility (Gorwa and Veale, 2024).
@@
Distribution platforms can amplify misuse by hosting and distributing models designed for AIG-NCII. Many early image-generation tools for generating AIG-NCII were hosted and distributed on GitHub, one of the most popular developer platforms in the world. In 2023, the National Center on Sexual Exploitation reported that GitHub had been hosting open-source repositories that enable AIG-NCII generation. This included DeepFaceLab, which “contains direct links to the most prolific sexual deepfake abuse website in the United States: Mr.DeepFakes,” as well as repositories for DeepNude and Unstable Diffusion (Patrick Trueman, Esq. and Dawn Hawkins, 2023).
@@
In the past several years, generative AI model distribution has increasingly shifted to CivitAI and Hugging Face. In May 2025, following new regulations and pressure from payment processors, CivitAI banned models designed to recreate real individuals’ likenesses, which, according to 404 Media, dealt a “major blow to the nonconsensual AI porn ecosystem” (Maiberg, 2025a, b). However, the deepfake community subsequently organized a coordinated effort to archive over 5,000 of these models and re-upload them to Hugging Face (Maiberg, 2025d). Recently, Hugging Face has been found to host models specifically designed for producing AIG-NCII of celebrities, including individuals who were minors when source material was created (Hashim, 2025; Maiberg, 2025d). These models violated Hugging Face’s stated content policy prohibiting “sexual content created without explicit consent” and “underage nudity or any sexual content involving minors” (Hugging Face, 2025). Nonetheless, Maiberg (2025d) found that Hugging Face was slow to take these models down, allowing some to remain accessible for nearly five months despite multiple requests. Even after removal events, some violating models remained accessible (Hashim, 2025).
Additionally, we searched across major model distribution platforms to investigate the scale of this moderation challenge. Searches for “dog whistle” terms commonly associated with AIG-NCII and AIG-CSAM production, including ‘realistic,’ ‘teen,’ ‘young,’ ‘girl,’ and ‘petite,’ returned over 26,000 models across the prominent platforms CivitAI and Hugging Face (Figure 4). [^12] Search methodology varied by platform capabilities. For CivitAI, we filtered for NSFW image and video generators; for Hugging Face, we filtered for image and video generators. We excluded Tensor.art and PixAI from this analysis because Tensor.art actively moderates NSFW content and PixAI does not have search counts. One of these terms appearing in a model’s name is not a clear sign that it specializes in photorealistic AIG-NCII/AIG-CSAM. However, this analysis suggests that platforms may host a substantial number of models designed for and capable of AIG-NCII/AIG-CSAM generation, particularly CivitAI.
Figure 4. Model search results across major distribution platforms for “dog whistle” terms often associated with AIG-NCII (January 2026). One of these terms appearing in a model’s name is not a clear sign that it specializes in photorealistic AIG-NCII/AIG-CSAM. However, this analysis suggests that platforms may host a substantial number of models designed for and capable of AIG-NCII/AIG-CSAM generation, particularly CivitAI.
This pattern illustrates how a lack of proactive enforcement of anti-AIG-NCII policies on model distribution platforms can make harmful models significantly more widely accessible.
@@ 5.1. Are AIG-NCII Harms from Open Video Generators Unmitigatable?
The “cat-out-of-the-bag” argument overlooks harm reduction principles and incorrectly characterizes technology adoption dynamics. The ‘cat-out-of-the-bag’ reasoning is not in line with established approaches to risk management and misidentifies how technology proliferates. Mitigations matter, especially in a domain where a substantial fraction of perpetrators are teenagers. For example, digital piracy provides one clear parallel: despite the theoretical availability of pirated content online, coordinated enforcement efforts by search engines, authorities, and companies to reduce the accessibility of pirated content online have had a measurable impact. For example, Smith (2023) concluded that removing links to pirated content from search results increases legal content sales by 11-14%, and (Danaher, 2017) concluded that search engines that demote piracy sites in results shift user behavior toward legal consumption. Even though determined users can still find pirated content with sufficient effort, increasing barriers to accessing it meaningfully reduces piracy by increasing search costs and directing casual users toward legitimate alternatives. Reducing the accessibility of harmful generative AI capabilities may be particularly helpful for reducing AIG-NCII given that a large number of perpetrators are minors themselves (Alexander, 2025).
Risk mitigation operates through aggregate effects, creating systemic benefits even when circumvention remains possible. Safety measures implemented by major platforms and developers create systemic benefits, including reducing total explicit content volume by avoiding accidental harms, preventing unsophisticated misuse, deterring costly circumventions of safeguards, preventing normalization by keeping harmful applications marginalized rather than mainstream, and enabling legal frameworks that can target deliberate safety circumvention (European Commission, 2019).
Technology adoption follows network effects that amplify early choices made by major actors. The first widely available implementations establish user expectations, development priorities, and community norms that influence subsequent development. For example, the LAION dataset was introduced in March 2022 (Beaumont and the LAION Team, 2022), was found to contain CSAM in December 2023 (Thiel, 2023), and was modified to mitigate CSAM content in August 2024 (LAION, 2024). Nonetheless, prior to August 2024, the original LAION dataset was widely used to train generative systems, including Stable Diffusion models (Salvaggio, 2024), that are still commonly used today. When major developers and platforms implement robust safety measures, they also create incentives for compatible approaches across the ecosystem. Conversely, early releases without safety measures normalize applications and create communities focused on circumvention rather than beneficial use. As such, the “cat-out-of-the-bag” argument offers a pretext to engage in a race to the bottom.